OpenStudy (anonymous):

I didn't understand this one. Let g(x) = 2x2 - 11x. Find g(-1).

OpenStudy (anonymous):

g (x) = 2x2 - 11x g (-1 ) = 2 ( -1) ^2 - 11 ( -1) you just need to replace the x by -1 and obtain the answer...

OpenStudy (anonymous):

just like y = 2x2 -11x and its telling you when x = -1 find y what you normally do is replace -1 in the equation y = 2x2 -11 x = 2 (-1) ^2 -11 (-11) :D

OpenStudy (anonymous):

I still don't understand. :C I'll post the answer choices.

OpenStudy (anonymous):

hep check your calculations ... g ( -1 ) = 2 (-1)^2 - 11 (-1) = 13

OpenStudy (anonymous):

2(-1)^2 = 2 and -11 (-1 ) = 11 2+11 = 13... just try again with your calculation the ans is D

OpenStudy (anonymous):

Okay.